The rate constant for the second-order reaction: 2NOBr(g) → 2NO(g) + Br2(g) is 0.80/(M · s) at 10°C. Starting with a concentration of 0.86 M, calculate the concentration of NOBr after 89 s. M
Answer #1

K= ţ ala-x) 9. initial concentration 0-86 M a = concentration of the reactant changed in time = !? (t = 895) (9-x) = concentr